OpenStudy (ambermarie151):

Solve this nonlinear equations. Show your work. Y=-x^2+6x-5 Y=3 A. Use the substitutions to combine equations. Rewrite so that one side is equal to 0. B. Factor your equation from Part 1. C. Identify the x-values of the solutions D. Identify the solutions to the system.

OpenStudy (jackthegreatest):

3 = x^2 + 6x - 5 0 = x^2 + 6x - 8 ...unless you meant -6x, this equation is unfactorable

OpenStudy (ambermarie151):

On the paper it says +6x @jackthegreatest

OpenStudy (jackthegreatest):

ahh i didnt see the - x^2

OpenStudy (jackthegreatest):

sorry....so: 0 = -x^2 + 6x - 8 0 = (-x +2)(x-4)

OpenStudy (jackthegreatest):

x = 2 or x = 4

OpenStudy (jackthegreatest):

@ambermarie151